# Optimal Control of Nonlinear Systems with Unknown Dynamics

### v2

@@ 6.2 Proof of Corollary 1
6.3 Proof of Lemma 2
To begin, we recall $\hat{J}_{\mu,t^{*}}$ from (17). For simplicity and ease of notation, we denote $\boldsymbol{u}_{t^{*}}^{i}\coloneqq\mu(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{i}^{\mu})$ and $\boldsymbol{\hat{x}}_{t^{*}+1}^{i}\coloneqq C_{k}(A_{k}\boldsymbol{g}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{k+1}^{f})+B_{k}\boldsymbol{u}_{t^{*}}^{i})$. By following the definition of $\hat{J}_{\mu,t^{*}}$, for any $\boldsymbol{\theta}_{1}^{\mu},\boldsymbol{\theta}_{2}^{\mu}\in\mathbb{R}^{q}$, the following holds:
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$ (31)
@@ 6.3 Proof of Lemma 2
×$\times\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{2})\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})-\gamma\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2})$
×$\times\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
@@
×$\times(\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{1}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu}))$
@@
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
@@
×$\times\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel+\gamma\parallel\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})-\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2})\parallel$
×$\times\parallel\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel.$
@@
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
@@
×$\times c_{2}c_{3}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel+\gamma L_{vxx}L_{\mu\theta}^{2}(c_{2}c_{3})^{2}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel$
@@
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}\parallel\leq(L_{c\mu}L_{\mu\theta}+\gamma L_{vx}L_{\mu\theta}c_{2}c_{3})\eqqcolon L_{J\theta}.$
[…]
6.4 Proof of Theorem 1
Recall $\hat{J}_{\mu}=\frac{1}{N}\sum_{t^{*}\in\mathcal{I}_{D}}\hat{J}_{\mu,t^{*}}$ from (15) as the averaged $\hat{J}_{\mu,t^{*}}$ in (17) computed over all $\boldsymbol{x}_{t^{*}}$ in $\mathcal{D}$. The goal of this proof is to analyze the convergence of $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}$ while tuning $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$ according to Algorithm 1.
We begin the analysis by employing the Taylor expansion for $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$ while disregarding higher-order terms, which will lead to the following results:
$\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$ (32)
$=$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})+\langle\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{\mu}-\boldsymbol{\theta}_{k}^{\mu}\rangle$
$+\frac{1}{2}\nabla_{\boldsymbol{\theta}^{\mu}\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel\boldsymbol{\theta}_{k+1}^{\mu}-\boldsymbol{\theta}_{k}^{\mu}\parallel^{2}$
$=$ $-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})^{\prime}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$+\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})+\frac{(\alpha_{k}^{\mu})^{2}}{2}\nabla_{\boldsymbol{\theta}^{\mu}\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$
×$\times\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}.$
Building on Lemma 2 and rearranging the results yields the following expression:
$\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$ (33)
$\leq$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$
@@ 6.4 Proof of Theorem 1
$=$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})+\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$-\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})+\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}.$
@@
$\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (34)
$\leq$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$
$+\langle\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}_{k}^{f}\rangle$
$+\langle\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}_{k}^{J}\rangle.$
@@
Here, to simplify the notation, we denote $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})=\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$, $\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f})=\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$ and $\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{J})=\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$ in the rest of this proof. Then taking the summation over the iteration steps $k=0,1,2,\cdots,K-1$ on both sides of (34) and applying Cauchy–Schwarz inequality yields:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (35)
$\leq$ $\sum_{k=0}^{K-1}\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}_{k}^{f}\parallel\parallel\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f})\parallel+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}_{k}^{J}\parallel$
×$\times\parallel\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{J})\parallel+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{K}^{f},\boldsymbol{\theta}_{K}^{J},\boldsymbol{\theta}_{K}^{\mu})$
$=$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})$
$-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}_{K}^{\mu})\!+\!\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}_{K}^{\mu})\!-\!\hat{J}_{\mu}(\boldsymbol{\theta}_{K}^{f},\boldsymbol{\theta}_{K}^{J},\boldsymbol{\theta}_{K}^{\mu})$
$+\sum_{k=0}^{K-1}\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}+\boldsymbol{\theta}^{f*}-\boldsymbol{\theta}_{k}^{f}\parallel\parallel\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f})\parallel$
$+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}^{J*}+\boldsymbol{\theta}^{J*}-\boldsymbol{\theta}_{k}^{J}\parallel\parallel\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{J})\parallel.$
Let $C_{1}=\max\{L_{vx}c_{1}c_{3},1\}$. Following Definition 1, triangle inequality, and Lipschitz continuity, we obtain the following bound:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (36)

### v3

@@ 6.1 Proof of Lemma 1
6.2 Proof of Theorem 1
Recall $\hat{J}=\frac{1}{N}\sum_{t^{*}\in\mathcal{I}_{D}}\hat{J}_{t^{*}}$ from (15) as the averaged $\hat{J}_{t^{*}}$ in (17) computed over $\mathcal{D}$. The goal of this proof is to analyze the convergence of $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}$ while tuning $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$ using the proposed PGDK framework.
6.2.1 Derivation of Lipschitz Constants
To begin, we need to ensure the approximated policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}$ from (16) is Lipschitz continuous. For simplicity and ease of notation, we denote $\boldsymbol{u}_{t^{*}}^{i}\coloneqq\mu(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{i}^{\mu})$ and $\boldsymbol{\hat{x}}_{t^{*}+1}^{i}\coloneqq C_{k}(A_{k}\boldsymbol{g}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{k+1}^{f})+B_{k}\boldsymbol{u}_{t^{*}}^{i})$. By following the definition of $\hat{J}_{t^{*}}$, for any $\boldsymbol{\theta}_{1}^{\mu},\boldsymbol{\theta}_{2}^{\mu}\in\mathbb{R}^{q}$, the following holds:
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$ (30)
@@ 6.2.1 Derivation of Lipschitz Constants
$\times\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{2})\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})-\gamma\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2})$
$\times\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
@@
$\times(\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{1}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu}))$
@@
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
@@
$\times\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel+\gamma\parallel\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})-\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2})\parallel$
$\times\parallel\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel.$
@@
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
@@
$\times c_{2}c_{3}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel+\gamma L_{vxx}L_{\mu\theta}^{2}(c_{2}c_{3})^{2}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel$
@@
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{t^{*}}\parallel\leq(L_{c\mu}L_{\mu\theta}+\gamma L_{vx}L_{\mu\theta}c_{2}c_{3})\eqqcolon L_{J\theta}.$
6.2.2 Convergence Analysis of the Estimated Policy Gradient
We now begin the analysis by employing the Taylor expansion for $\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$ while disregarding higher-order terms, which will lead to the following results:
$\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$ (31)
$=$ $\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})+\langle\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{\mu}-\boldsymbol{\theta}_{k}^{\mu}\rangle$
$+\frac{1}{2}\nabla_{\boldsymbol{\theta}^{\mu}\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel\boldsymbol{\theta}_{k+1}^{\mu}-\boldsymbol{\theta}_{k}^{\mu}\parallel^{2}$
$=$ $-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})^{\prime}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$+\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})+\frac{(\alpha_{k}^{\mu})^{2}}{2}\nabla_{\boldsymbol{\theta}^{\mu}\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$\times\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}.$
Since the approximated policy gradient is Lipschitz continuous and rearranging the results yields the following expression:
$\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$ (32)
$\leq$ $\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$
@@ 6.2.2 Convergence Analysis of the Estimated Policy Gradient
$=$ $\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})+\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$-\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})+\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}.$
@@
$\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (33)
$\leq$ $\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$
$+\langle\nabla_{\boldsymbol{\theta}^{f}}\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}_{k}^{f}\rangle$
$+\langle\nabla_{\boldsymbol{\theta}^{J}}\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}_{k}^{J}\rangle.$
@@
Here, to simplify the notation, we denote $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})=\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$, $\nabla_{\boldsymbol{\theta}^{f}}\hat{J}(\boldsymbol{\theta}_{k}^{f})=\nabla_{\boldsymbol{\theta}^{f}}\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$ and $\nabla_{\boldsymbol{\theta}^{J}}\hat{J}(\boldsymbol{\theta}_{k}^{J})=\nabla_{\boldsymbol{\theta}^{J}}\hat{J}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$ in the rest of this proof. Then taking the summation over the iteration steps $k=0,1,2,\cdots,K-1$ on both sides of (33) and applying Cauchy–Schwarz inequality yields:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (34)
$\leq$ $\sum_{k=0}^{K-1}\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}_{k}^{f}\parallel\parallel\nabla_{\boldsymbol{\theta}^{f}}\hat{J}(\boldsymbol{\theta}_{k}^{f})\parallel+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}_{k}^{J}\parallel$
$\times\parallel\nabla_{\boldsymbol{\theta}^{J}}\hat{J}(\boldsymbol{\theta}_{k}^{J})\parallel+\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}(\boldsymbol{\theta}_{K}^{f},\boldsymbol{\theta}_{K}^{J},\boldsymbol{\theta}_{K}^{\mu})$
$=$ $\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})$
$-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}_{K}^{\mu})\!+\!\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}_{K}^{\mu})\!-\!\hat{J}(\boldsymbol{\theta}_{K}^{f},\boldsymbol{\theta}_{K}^{J},\boldsymbol{\theta}_{K}^{\mu})$
$+\sum_{k=0}^{K-1}\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}+\boldsymbol{\theta}^{f*}-\boldsymbol{\theta}_{k}^{f}\parallel\parallel\nabla_{\boldsymbol{\theta}^{f}}\hat{J}(\boldsymbol{\theta}_{k}^{f})\parallel$
$+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}^{J*}+\boldsymbol{\theta}^{J*}-\boldsymbol{\theta}_{k}^{J}\parallel\parallel\nabla_{\boldsymbol{\theta}^{J}}\hat{J}(\boldsymbol{\theta}_{k}^{J})\parallel.$
Let $C_{1}=\max\{L_{vx}c_{1}c_{3},1\}$. Following the triangle inequality, and Lipschitz continuity, we obtain the following bound:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (35)